Arsenic can be determined by a redox titration with standard iodine solution. The reaction is HAsO2 + I2 → H3AsO4 + 2H+ + 2IPure As2O3 can be used as a primary standard in determining the concentration (normality) of the iodine solution. A sample of pure As2O3 weighing 4.0136 g is dissolved in 800.0 mL of solution. Calculate the normality of the solution when it is used in the reaction above. Also calculate the molarity of the solution. Note that each arsenic atom loses two electrons in being oxidized from HAsO2 to H3AsO4. Each As2O3 contains two arsenic atoms and hence loses four electrons during the standardization reaction: As2O3 → 2 HAsO2 → 2 H3AsO4 + 4eTherefore, the equivalent weight of As2O3 is ____ of the molecular weight.).

PLEASE PUT EXPLANATION IN THE ANSWERS.
1
Expert's answer
2021-10-11T06:12:16-0400

As2O3 = 4.0136 g

V = 800 mL

a)

normality of solution when used...

first, get moles

MW of AsO3 = 122.91980 g/mol

mol = mass/MW 4.0136/122.91980 = 0.03265 mol

now..

[AsO3] = mol/V = 0.03265 / 0.8 = 0.04081 M

now...

As2O3 → 2 HAsO2 → 2 H3AsO4

Normality -- > 0.04081 M * 2 equivalents = 0.08162 equivalents / Liter = 0.08162 Normality